# Explore Spurious Correlations at the Concept Level in Language Models for Text Classification

### 06 Jan 2024 v2

@@ 1 Introduction
Pre-trained language models (LMs), leveraging extensive text corpora in their pre-training phase, have demonstrated remarkable effectiveness in a variety of natural language understanding tasks Wei et al. (2022); Devlin et al. (2018). Nevertheless, LMs encounter issues with spurious correlations during fine-tuning or instruction-following stages Zhang et al. (2022); Wang et al. (2022); Tang et al. (2023). These correlations involve specific associations between features and labels that, while prevalent in training data, are erroneously generalized as rules, leading to reduced performance in downstream tasks.
@@ 2.1 Obtaining the Concept Labels
Due to the lack of human-annotated metadata indicating concepts in most existing text classification datasets, and considering the superior capabilities of LLMs in text annotation tasks over human annotators Gilardi et al. (2023), we utilize ChatGPT (GPT-3.5) to annotate concept labels for sentences in text classification datasets Ouyang et al. (2022). Our annotation process involves an annotation prompt $P_{a}$ that contains the annotation instruction and five demonstrations, a text input $x$, LLM $M_{a}$, and a candidate concept set $C=\{C_{1},C_{2},\cdots,C_{k}\}$ (with dataset-specific details provided in Section 3).
@@ 2.2 Measuring Concept Spurious Correlations
For the text classification task, we consider a text input $x\in\mathcal{X}$ accompanied by concept labels $a(x)\subset C$. Each input is associated with a ground truth classification label $y=l$ from the output label space $\mathcal{Y}$, $l\in\{0,1,\cdots,n\}$. Given a LM classifier $M:\mathcal{X}\rightarrow\mathcal{Y}$, if the model avoids utilizing potential concept-level shortcuts from $c\rightarrow y,c\in C$, the following condition is satisfied:
$\mathbb{E}_{x}[p_{M}(\hat{y}=0|x,c\in a(x),y=l)]$ (1)
$=$ $\mathbb{E}_{x^{\prime}}[p_{M}(\hat{y}=0|x^{\prime},c\notin a(x^{\prime}),y=l)]\quad\forall l\in\mathcal{Y}.$
Here, $\hat{y}$ denotes the predicted label, while $p_{M}$ represents the probability predicted by model $M$. The inputs $x$ and $x^{\prime}$, belonging to the space $\mathcal{X}$, contain the concept $c$ or do not contain it, respectively. Equation 1 implies a critical condition: irrespective of the presence of concept $c$ in the input, the models should maintain an unbiased estimate of the predicted probability on average. The expression $\mathbb{E}_{x}[p_{M}(\hat{y}=l|x,c\in a(x),y=l)]$ can be interpreted as the model’s accuracy on texts that are labeled $l$ and incorporate the concept $c$.
Equation 1 allows us to infer that the difference in model accuracy between any two classification labels, $i$ and $j$, should be independent of the concept label $c$. Define the accuracy difference for texts containing concept $c$ between labels $i$ and $j$ as $\Delta_{c_{ij}}$, and for texts lacking concept $c$ as $\Delta_{\neg c_{ij}}$. This leads to the following formulation:
$\begin{split}&\Delta_{c_{ij}}=\mathbb{E}_{x}[p_{M}(\hat{y}=i|x,c\in a(x),y=i)]-\mathbb{E}_{x}[p_{M}(\hat{y}=j|x,c\in a(x),y=j)]\\ &\Delta_{\neg c_{ij}}=\mathbb{E}_{x^{\prime}}[p_{M}(\hat{y}=i|x^{\prime},c\notin a(x^{\prime}),y=i)]-\mathbb{E}_{x^{\prime}}[p_{M}(\hat{y}=j|x^{\prime},c\notin a(x^{\prime}),y=j)].\\ \end{split}$
Consequently, we arrive at the equality $\Delta_{c_{ij}}=\Delta_{\neg c_{ij}}$, allowing us to quantify the model’s reliance on shortcut mapping as the average discrepancy in accuracy difference, $\Delta_{c_{ij}}-\Delta_{\neg c_{ij}}$, across all label combinations for inputs with and without $c$.
@@
2.3 Evaluation of Shortcut Robustness in Fine-tuning
To assess LMs’ robustness against spurious correlations for concept $c$ across varying scales of concept bias during fine-tuning, we fine-tune models on both the original dataset $\mathcal{D}_{ori}$ and a concept-biased dataset $\mathcal{D}^{c}_{biased}$. Dataset $\mathcal{D}^{c}_{biased}$ is constructed by filtering $\mathcal{D}_{ori}$. If the majority classification labels for sentences under concept $c$ are $\{l_{1},l_{2},\cdots,l_{k}\}$, where $k=\lceil\frac{n}{2}\rceil$, we will preserve the majority labels to enforce $p(y\in\{l_{1},l_{2},\cdots,l_{k}\}|c\in a(x))=1$ in $\mathcal{D}^{c}_{biased}$. In binary classification, $\mathcal{D}^{c}_{biased}$ retains only texts with concept $c$ labeled exclusively as 0 or 1. Post fine-tuning on $\mathcal{D}_{ori}$ or $\mathcal{D}^{c}_{biased}$, we evaluate the models on test data using Bias@C to quantify spurious correlations.
To measure the utility performance of fine-tuned models, we calculate the accuracy results on the test data, but the label distribution for the text in the test with or without the concept $c$ may not be balanced. Following the previous work Chew et al. (2023), we rebalance the label by downsampling the examples and report the inference accuracy (robust accuracy) on the balanced subset for texts with concept $c$ (Acc@C) and without concept $c$ (Acc@NoC), respectively.
2.4 Evaluation of Shortcut Robustness in ICL
After measuring concept-level spurious correlations in model fine-tuning, we ask whether the LMs use the shortcut mapping in the prompt to perform the ICL. Our prompt $P$ for ICL contains three parts: 1) the instruction $s$ to introduce the information of the task, 2) the demonstrations with $h$ exemplars (text + labels), and 3) the test input $x_{test}$.
@@ 2.4 Evaluation of Shortcut Robustness in ICL
We create two types of prompts: the biased prompt $P_{biased}$ and the balanced prompt $P_{balanced}$ by changing the label distributions in the demonstrations. For $P_{biased}$, we insert $\frac{h}{2}$ numbers of exemplars containing the concept $c$ with the label $l\in\{l_{1},l_{2},\cdots,l_{k}\}$ (the majority ground truth labels) and $\frac{h}{2}$ numbers of exemplars not containing $c$ with the other labels. For $P_{balanced}$, we still split the exemplars with concept $c$ or not half by half, but we ensure balanced labels. We compare the results of Bias@C and robust accuracy with two types of prompts. Since the ICL inference are sensitive to the exemplars, we repeat the experiments three times with differently selected exemplars and report the average values.
@@ 3 Dataset Construction and Analysis
Dataset  We select four sentiment classification tasks to evaluate the model robustness: Yelp Zhang et al. (2015), IMDB Maas et al. (2011), Amazon Shoe He and McAuley (2016), and CeBaB Abraham et al. (2022). Amazon shoe and CeBaB datasets with 5 classes, 0 indicating the most negative and 4 indicating the most positive, are reviews of shoes in Amazon and OpenTable. IMDB and Yelp are binary classification datasets (0 indicating negative and 1 indicating positive), with reviews from the IMDB and Yelp platforms. In the following part, we define the positive texts as texts with Label 3 and 4 for the 5-class classification task and Label 1 for the binary classification task. We define the remaining texts as the negative texts. Note that the concept labels of Yelp, IMDB, and Amazon shoe are annotated by ChatGPT, and for CeBaB, the labels are from human annotators.
Concept  For the concept list for Amazon Shoe, IMDB, and Yelp without the meta concept annotation, we first use ChatGPT to query the concepts embedded in each sentence and count the number of occurrences for each concept following Fang and Zhang (2022) to generate concept-level explanations. We extract the top concepts and identify the concepts whose existence should not influence the sentiments of the text. Then, we use ChatGPT to annotate the text that contains the concept or not. To examine the precision of the annotation, we experiment on the human-annotated “service” concept from the CeBaB dataset and ask the ChatGPT to label the concept again. We find that ChatGPT can achieve an accuracy of 90.4% to the gold standard concept labels, comparable to an average accuracy of 92.9% for five human annotators given by CeBaB, indicating that the reliability of LLM annotations. In Table 1, we present dataset statistics and the labeled concept lists for the four datasets.
@@ 3.1 Biased Dataset Construction
Given that the overall label distribution is nearly balanced for each dataset, our first exploration is to visualize the label distribution for the input texts with the concept $c$ for each dataset in Figure 3. We observe that for the original dataset, the label distributions are not balanced for concepts such as “food” in Yelp dataset, “music” in IMDB dataset, and “style” in Amazon Shoe dataset. In most cases, positive labels comprise large proportions of the corpus with certain concepts, except for “size” in Amazon Shoe dataset and “service” in Yelp dataset. Therefore, to construct the biased dataset $\mathcal{D}^{c}_{biased}$, we keep negative texts for “size” in Amazon Shoe and “service” in Yelp and for other concepts, we keep positive texts.
@@ 3.2 Embedding Analysis of Associated Tokens
From the associated tokens in Table 2, we observe various types of tokens of one concept, such as “small,” “sizing” and “9m” to express the “size” concept.
We use the tokens in Table 2 to perform the clustering. We exclude tokens with special character, such as “c/d,” “81/2” and “(55)” that affects the interpretation of the results. We feed the tokens into the DistilBERT fine-tuned on the Amazon Shoe dataset and retrieve the corresponding embedding from the model last layer. If the token is tokenized into multiple sub-words, we follow the previous work and calculate the average as the token embedding Wolfe and Caliskan (2021). We calculate the cosine distance between their token embeddings and apply hierarchical clustering Bar-Joseph et al. (2001).
From the visualization in Figure 4, we can identify four small clusters, each representing a concept. We observe that the LMs will produce similar internal representations for tokens associated with the same concept label. If the label under a concept is imbalanced, the models may learn the undesired shortcut between similar embeddings and a label. This observation motivates the measurement of spurious correlation at the concept level.
@@ 4.1 Spurious Correlations in Fine-tuning
To evaluate the robustness of the shortcut in the fine-tuning setting, we first fine-tune the models on the original dataset $\mathcal{D}_{ori}$ and the biased dataset $\mathcal{D}^{c}_{biased}$, respectively, and measure the concept bias. For each concept, we report the metric Bias@C to quantify the strength of spurious correlations and the robust accuracy for texts with and without concept, i.e., Acc@C and Acc@NoC, as the utility performance of the fine-tuned model. For Bias@C, closer to 0 indicates weaker spurious correlations for concept $C$, and for robust accuracy, a higher value suggests better performance. We present the results for DistilBERT in Table 3 and discuss the findings as follows.
Fine-tuned LMs present a clear concept bias when trained on both original and biased data.  Table 3 shows that when the models are trained on $\mathcal{D}_{ori}$, they utilize spurious correlations presented in the training data to make inferences. For example, for “style” in the Amazon Shoe and “music” in IMDB, Bias@C values for the fine-tuned models in $\mathcal{D}_{ori}$ are large.
Comparing the results between $\mathcal{D}_{ori}$ and $\mathcal{D}^{c}_{biased}$, we find that the absolute values of Bias@C are significantly higher when trained on $\mathcal{D}^{c}_{biased}$ in almost every concept, and the change direction of Bias@C is the same as the trend in the label distribution. For example, the value of Bias@C becomes negative for “service” in Yelp dataset, since we only keep negative reviews with the “service” concept in $\mathcal{D}^{c}_{biased}$. These observations indicate that a greater bias in the fine-tuning dataset causes the model to rely more on spurious correlations to make predictions.
Regarding utility performance (Acc@C and Acc@NoC), we observe that the models trained on $\mathcal{D}^{c}_{biased}$ have a dramatic performance drop on the texts with the concept in most cases, and the average Acc@C decreases from 79.38% to 74.31%. This pattern suggests that larger spurious correlations affect the utility performance of fine-tuned models. However, for Acc@NoC, this performance drop is not as large as the change in Acc@C, and the average Acc@NoC drops from 78.08% to 76.56%, which means that texts without the concept suffer less influence. We also verify that the concept-level bias is not simply due to the shortcut on a few words by masking out the associated tokens, and details are shown in Section 4.1. We repeat the experiments on LLAMA2 7B with four datasets and present the results in Table 6 in Appendix B. Similar patterns suggest the generalizability of our findings on models of different sizes.
@@
Table 4: Model ICL performance with prompting on balanced prompts $P_{balanced}$ and biased prompts $P_{biased}$ for four datasets. Larger absolute values for Bias@C indicate that the concept-biased prompts can enlarge the extend of models to rely on the shortcut in the demonstrations. The meaning of “pos > neg” is the same as in Table 3. The lower absolute values of Bias@C (smaller bias) and the higher values of accuracy are in bold.

### 21 Feb 2024 v3

@@ 1 Introduction
Pre-trained language models (LMs), leveraging extensive text corpora in their pre-training phase, have demonstrated remarkable effectiveness in a variety of natural language understanding tasks Wei et al. (2022); Devlin et al. (2018). Nevertheless, LMs encounter issues with spurious correlations during fine-tuning or instruction-following stages Zhang et al. (2022); Wang et al. (2022); Tang et al. (2023). These correlations involve specific associations between features and labels that, while prevalent in training data, are erroneously generalized as rules, leading to reduced performance.
@@ 2.1 Obtaining the Concept Labels
Due to the lack of human-annotated metadata indicating concepts in most text classification datasets, and considering the superior capabilities of LLMs in text annotation tasks over human annotators Gilardi et al. (2023), we utilize ChatGPT (GPT-3.5) to annotate concept labels for sentences in text classification datasets Ouyang et al. (2022). Our annotation process involves an annotation prompt $P_{a}$ that contains the annotation instruction and five demonstrations, a text input $x$, LLM $M_{a}$, and a candidate concept set $C=\{C_{1},C_{2},\cdots,C_{k}\}$ (we describe how we curate the candidate set in Section 3).
@@ 2.2 Measuring Concept Spurious Correlations
For the text classification task, we consider an input $x\in\mathcal{X}$ accompanied by concept labels $a(x)\subset C$. Each input is associated with a ground truth classification label $y=l$ from the output label space $\mathcal{Y}$, $l\in\{0,1,\cdots,n\}$. Given a LM classifier $M:\mathcal{X}\rightarrow\mathcal{Y}$, if the model avoids utilizing potential concept-level shortcuts from $c\rightarrow y,c\in C$, the following condition is satisfied:
$\mathbb{E}_{x}[p_{M}(\hat{y}=l|x,c\in a(x),y=l)]$ (1)
$=$ $\mathbb{E}_{x^{\prime}}[p_{M}(\hat{y}=l|x^{\prime},c\notin a(x^{\prime}),y=l)]\quad\forall l\in\mathcal{Y}.$
Here, $\hat{y}$ denotes the predicted label, while $p_{M}$ represents the probability predicted by model $M$. The inputs $x$ and $x^{\prime}$, belonging to the space $\mathcal{X}$, contain the concept $c$ or do not contain it, respectively.
Equation 1 implies a critical condition: regardless of the presence of concept $c$ in the input, the models should maintain an unbiased estimate of the predicted probability on average. The expression $\mathbb{E}_{x}[p_{M}(\hat{y}=l|x,c\in a(x),y=l)]$ can be interpreted as the model’s accuracy on texts that are labeled $l$ and incorporate the concept $c$.
Denote $\Delta_{c_{i}}$ as the difference in model accuracy between texts with or without concept $c$ that have label $i\in\mathcal{Y}$. We further infer from Equation 1 that:
$\Delta_{c_{i}}=$ $\mathbb{E}_{x}[p_{M}(\hat{y}=i|x,c,y=i)]$ (2)
-- $\mathbb{E}_{x^{\prime}}[p_{M}(\hat{y}=i|x^{\prime},\neg c,y=i)]=0,$
where $\neg c$ denotes concept $c$ is not in input $x$. We hypothesize, if there exists a spurious correlation in models between concept $c$ and label $i$, the following conditions would hold:
[…]
Then we have $\Delta_{c_{i}}>0>\Delta_{c_{j}}$. Otherwise, if the spurious correlation is between $c$ and $j$, then $\Delta_{c_{j}}>0>\Delta_{c_{i}}$. We propose to measure the discrepancy between $\Delta_{c_{i}}$ and $\Delta_{c_{j}}$ to quantify the spurious correlation. Hence, considering the output space $\mathcal{Y}$, we quantify the model’s reliance on shortcut mapping as the average discrepancy in the accuracy difference $\Delta_{c_{i}}-\Delta_{c_{j}}$ across all label combinations.
@@
A Bias@C approaching 0 indicates minimal reliance on concept shortcuts. Conversely, a positive Bias@C value suggests that model is more likely to predict larger labels when the input includes concept $c$, and the opposite for a negative value.
2.3 Evaluation of Model Robustness to Concept Shortcut in Fine-tuning
To assess LMs’ robustness against spurious correlations for concept $c$ across varying scales of concept bias during fine-tuning, we fine-tune models on the original dataset $\mathcal{D}_{ori}$ and a concept-biased dataset $\mathcal{D}^{c}_{biased}$ separately. To further demonstrate the impact of concept-level spurious correlation, we construct $\mathcal{D}^{c}_{biased}$ of concept $c$ by filtering $\mathcal{D}_{ori}$, where, for each data point, we only keep those with the majority labels under concept $c$. After fine-tuning on $\mathcal{D}_{ori}$ or $\mathcal{D}^{c}_{biased}$, we evaluate models on test data using Bias@C to quantify spurious correlations.
We report accuracy on the test data for utility performance. However, label distributions with or without the concept $c$ may be imbalanced. Following previous work Chew et al. (2023), we rebalance the test set by downsampling and report the inference accuracy (robust accuracy) on the balanced subset for examples with concept $c$ (Acc@C) and without concept $c$ (Acc@NoC), respectively.
2.4 Evaluation of Model Robustness to Concept Shortcut in ICL
As LLMs have shown outstanding performances with the ICL setting, we are interested in investigating the concept shortcut in the demonstrations. The prompt $P$ for ICL contains three parts: 1) the instruction $s$, 2) the demonstrations with $h$ exemplars (text + labels), and 3) the test input $x_{test}$.
@@ 2.4 Evaluation of Model Robustness to Concept Shortcut in ICL
We create two types of prompts: the biased prompt $P_{biased}$ and the balanced prompt $P_{balanced}$ by changing the label distributions in the demonstrations. For $P_{biased}$, we insert $\frac{h}{2}$ numbers of exemplars containing the concept $c$ with the label $l\in\{l_{1},l_{2},\cdots,l_{k}\}$ (the majority ground truth labels) and $\frac{h}{2}$ numbers of exemplars without $c$ with the other labels. For $P_{balanced}$, we split the exemplars with concept $c$ or not half by half, but ensure balanced labels. We compare the results of Bias@C and robust accuracy with two types of prompts. Since the ICL are sensitive to the exemplars, we repeat the experiments three times with differently selected exemplars and report the average values.
@@ 3 Dataset Construction and Analysis
Dataset  We select four sentiment classification tasks to evaluate the model robustness: Yelp Zhang et al. (2015), IMDB Maas et al. (2011), Amazon Shoe He and McAuley (2016), and CeBaB Abraham et al. (2022). Amazon shoe and CeBaB datasets with 5 classes, 0 indicating the most negative and 4 indicating the most positive, are reviews of shoes in Amazon and OpenTable. IMDB and Yelp are binary classification datasets (0 indicating negative and 1 indicating positive), with reviews from the IMDB and Yelp platforms.
Additionally, we include a binary question answering (QA) dataset BoolQ Clark et al. (2019), which asks Yes/No questions. It takes a paired question and passage as the input to LMs and outputs 1 (Yes) or 0 (No). In the following part, we define the positive class as datapoints with Label 3 and 4 for the 5-way classification tasks and those with Label 1 for the binary classification task. We define the remaining datapoints as the negative class.
Concept  For CeBaB, we adopt human-annotated concept labels. For Amazon Shoe, IMDB, Yelp, and BoolQ where there are no concept annotations, we first use ChatGPT to query the concepts embedded in each sentence and count the number of occurrences for each concept following Fang and Zhang (2022) to generate concept-level explanations. We then extract the most frequent concepts and identify the concepts whose existence should not influence the sentiments of the text or the Yes/No answer to the question (2 concepts for BoolQ due to more diverse topics and 3 concepts for other datasets). Finally, we use ChatGPT to annotate whether each text input contains the selected concept.
To examine the quality of the annotation, we experiment on the human-annotated “service” concept from the CeBaB dataset and ask the ChatGPT to label the concept again. We find that ChatGPT can achieve an accuracy of 90.4% to the gold standard concept labels, comparable to an average agreement of 92.9% for five human annotators given by CeBaB, indicating the reliability of LLM annotations. Table 1, presents dataset statistics and the labeled concept lists for the five datasets.
@@ 3.1 Biased Dataset Construction
We first visualize the label distribution for the input texts with the concept $c$ for each sentiment classification in Figure 3. We observe that for the original datasets, the concept-label distributions are balanced for most concepts, but not as balanced for concepts such as “food” in Yelp dataset, “music” in IMDB, and “style” in Amazon Shoe. In 10/12 cases, positive labels comprise large proportions of the corpus with certain concepts. To further demonstrate the impact of concept-level spurious correlation caused by imbalanced concept-label distribution, we construct a biased dataset $\mathcal{D}^{c}_{biased}$ which, for each concept $c$, only includes the majority class (positive or negative). Specifically, we keep negative class for “size” in Amazon Shoe and “service” in Yelp. For other concepts in the sentiment datasets, we keep positive class. For BoolQ, we keep negative class with “country” and positive class with “history”.
@@ 3.2 Embedding Analysis of Associated Tokens
From the associated tokens in Table 2, we observe tokens with various semantics associated with one concept, such as “small,” “sizing” and “9m” to express the “size” concept. We use the tokens in Table 2 to perform the clustering. We exclude tokens with special character, such as “c/d,” “81/2” and “(55)” that affect the interpretation of the results. We feed the tokens into the DistilBERT fine-tuned on the Amazon Shoe and retrieve the corresponding embedding from the model last layer. If the token is tokenized into multiple sub-words, we follow the previous work and calculate the average as the token embedding Wolfe and Caliskan (2021). We calculate the cosine distance between their token embeddings and apply hierarchical clustering Bar-Joseph et al. (2001).
From Figure 4, we can identify four small clusters, each representing a concept. We observe that the LMs will produce similar internal representations for tokens associated with the same concept label. If the label under a concept is imbalanced, the models may learn the undesired shortcut between similar embeddings and a label. This observation motivates the measurement of spurious correlation at the concept level.
@@ 4.1 Spurious Correlations in Fine-tuning
To evaluate the robustness to the concept shortcut in the fine-tuning setting, we fine-tune the models on the original dataset $\mathcal{D}_{ori}$ and the biased dataset $\mathcal{D}^{c}_{biased}$, respectively, and measure the concept bias. For each concept, we report the metric Bias@C to quantify the strength of spurious correlations and the robust accuracy for texts with and without concept, i.e., Acc@C and Acc@NoC, as the utility performance. For Bias@C, closer to 0 indicates weaker spurious correlations for concept $C$, and for robust accuracy, a higher value suggests better performance. We present the results for DistilBERT on sentiment classification in Table 3 and BoolQ dataset in Table 6 Appendix B.
Fine-tuned LMs present a clear concept bias when trained on both original and biased data.  Table 3 shows that when the models are trained on $\mathcal{D}_{ori}$, they utilize spurious correlations in the training data to make inferences. For example, for “style” in the Amazon Shoe and “music” in IMDB, the Bias@C values in $\mathcal{D}_{ori}$ are large due to highly imbalanced label distribution.
Comparing the results between $\mathcal{D}_{ori}$ and $\mathcal{D}^{c}_{biased}$, we find that the absolute values of Bias@C are significantly higher when trained on $\mathcal{D}^{c}_{biased}$ in almost every concept, and the change direction of Bias@C is the same as the trend in the label distribution. For example, the value of Bias@C becomes negative for “service” in Yelp dataset, since we only keep negative reviews with the “service” in $\mathcal{D}^{c}_{biased}$. These observations indicate that a greater bias in the fine-tuning dataset causes the model to rely more on spurious correlations to make predictions.
Regarding utility performance (Acc@C and Acc@NoC), we observe that the models trained on $\mathcal{D}^{c}_{biased}$ have a dramatic performance drop on the texts with the concept in most cases, and the average Acc@C decreases from 79.38% to 74.31%. This pattern suggests that larger spurious correlations affect the utility performance of fine-tuned models. Meanwhile, the average Acc@NoC drops from 78.08% to 76.56%. Its performance drop is not as large as the one of Acc@C, indicating that texts without the concept suffer less from the concept bias in the datasets.
We also verify that the concept bias is not simply due to the shortcut on a few words by masking out the associated tokens, and details are shown in Section 5.2. We show fine-tuning results of LLAMA2 models on $\mathcal{D}_{ori}$ and $\mathcal{D}^{c}_{biased}$ in Table 6 and 7 (Appendix B). Similar patterns suggest the generalizability of our findings on models of different sizes.
@@
Table 4: Model ICL performance with prompting on balanced prompts $P_{balanced}$ and biased prompts $P_{biased}$. Larger absolute values for Bias@C indicate that the concept-biased prompts enlarge the extend of models to rely on the shortcut in the demonstrations. The meaning of “pos > neg” and the values in bold are the same as in Table 3.